a circular fountain 10 ft. in diameter has a walk 3 ft wide around it. what is the area of the walk?
Answer by jessica43(140)

You are going to need to find the area of the walk and the fountain together, and then subtract the area of the fountain by itself to find the area of just the walk.
First we will find the area of the fountain. Area of a circle= π*r^2, where r= radius. I am going to use 3.14 as an approximation for π.
Since the diameter is 10, then the radius of the fountain is 5, so the area is:
A = π* r^2
A = π(5^2)
A = 25π
A= 78.5
Now find the area of the fountain and the walk, so the diameter is now 13 ft, and the radius is 6.5 ft:
A =π*r^2
A = π(6.5^2)
A =42.25π
A = 132.665
Now subtract the area of the fountain from the area of the fountain and walk together. What is left is just the area of the walk.
132.665 - 78.5 = 54.165
So the area of the walk is 54.165 (or more specifically 42.25π - 25π = 17.25π)